# Applying the maximum entropy principle to neural networks enhances multi-species distribution models

### v3

@@ 2.1.3 Batched algorithm and partition function approximation in DeepMaxent
One of the challenges of adapting Maxent to a deep learning framework is the computation of the partition functions $\sum_{k}\lambda_{kj}$ for each species $j$, corresponding to the denominator in equation 2, which normalises the predicted intensity $\lambda_{j}(x)$ over the $K$ sites. When the geographic domain becomes large, or when the spatial resolution is increased, the number of terms $K$ becomes large, and computing this function may be particularly expensive in a deep learning context where each $g_{\theta}(x_{i})$ may already be costly. To address this consistently with the batched optimisation algorithms used in deep learning, such as Stochastic Gradient Descent (SGD), we modify the normalisation step by computing the loss function on a small random subset of sites $B\subset[1,K]$, called batch, hence normalising the intensities within the batch. The batch-wise loss is then written as follows:
@@
$\mathcal{L}(\tilde{\lambda}_{i\in B},\tilde{y}_{i\in B})=-\frac{1}{|B|N}\sum_{i\in B}\sum_{j=1}^{N}\frac{y_{ij}}{\sum_{i\in B}y_{ij}}\log\left(\frac{\lambda_{j}(x_{i})}{\sum_{i\in B}\lambda_{j}(x_{i})}\right).$ (7)
@@
This batch-wise loss makes the model optimisation computationally feasible by computing it one batch at a time, and thus allows it to be scalable to large domains and trained efficiently with any stochastic optimiser based on random batches, such as the very commonly used method Adam which we use here (Kingma (2014)). However, it must be noted that the batch-wise loss is not a simple approximation of the full loss, as for instance the normalised labels tend to have a larger value for smaller batches. Nevertheless, we provide the mathematical guarantee that, for any batch size $n$ ($1<n<K$), optimising the model on all batches also optimises the full loss (see Appendix A.3). This suggests that our final estimator should be somehow close to the global minimiser of the full loss, even though we will unlikely obtain the latter. Similarly to supervised contrastive methods Khosla et al. (2020), the intensity predictions could become more specific, and more concentrated around the occurrences, as the batch size increases. A small batch size could, on the other hand, result in smoother species intensities. Therefore we tested the impact of the batch size during training on the final predictions of DeepMaxent. Given the expression of $\lambda_{j}(x_{i})$ in equation 6, its normalisation across the batch $B$ writes as $\exp(\gamma_{j}^{T}g_{\theta}(x_{i}))/\sum_{k\in B}\exp(\gamma_{j}^{T}g_{\theta}(x_{k}))$. This is a particular case of a softmax function applied to the logits $\gamma_{j}^{T}g_{\theta}(x_{k})$ over the batch $B$.
@@ 2.1.4 Spatial sampling bias correction with Target-Group Background correction
When occurrence concentration is biased by spatial variations in sampling effort, a popular SDM correction approach is the Target-Group Background (TGB) method Phillips et al. (2009), which was initially proposed to correct sampling bias in Maxent. The method basically approximates the spatial sampling effort through the distribution of occurrences of a Target-Group of species, providing background points to Maxent for each site where TG species were reported. In other words, TGB restricts the study domain to the sites with at least an evidence of sampling effort (one observation), which reduces the problem of false absences. The strategy is expected to work when TG species are reported jointly with the focal species (e.g. the TG is a biological group targeted by a citizen science program).
In DeepMaxent, which is a multi-species model, it can be relevant to define the whole set of $N$ predicted species as the Target-Group. This is done in the empirical study below: For each region, each biological group is the set of predicted species and the Target-Group, so the sites where at least one of the these TG species was observed are the background sites. Then, for each iteration of our batched algorithm, we randomly select a same subset of TG background sites for all predicted species.
@@ 2.1.5 L2-regularisation implementation in DeepMaxent
Maxent makes use of a L1 penalisation, on the species weights $\gamma_{j}$ that model the relation between the features and the density prediction. The L1 term, known as LASSO penalty, encourages $\gamma_{j}$ to become sparse, thus selecting a subset of features.
The L1 regularisation is important in Maxent due to a number of features that grows more than quadratically with the number of environmental variables provided a priori by the user. In DeepMaxent, the latent features are learned to maximise prediction performances, removing the need for feature selection. For DeepMaxent, we employ a L2 regularisation, i.e. a penalty on the Euclidean norm of the $\gamma_{j}$, which encourages small but non-zero weights, which may induce a smoothing of the estimated species intensities. The total loss function thus becomes:
@@ 2.2 Evaluation of model performance
2.2.1 Dataset
For our experiments we used an openly released dataset from the National Centre for Ecological Analysis and Synthesis (NCEAS) (Elith et al., 2020). This dataset includes PO (for SDM training) and presence-absence (PA, for SDM evaluation) data from six global regions: Australian Wet Tropics (AWT), Canada (CAN), New South Wales (NSW), New Zealand (NZ), South America (SA) and Switzerland (SWI) (Elith et al., 2020). Each regions is associated with a specific set of species, and sometimes from several biological groups (see Table 1), with a total of 226 anonymous species. The dataset provides specific environmental variables for each region, including climatic, soil or location variables (see more details Elith et al. (2020)).
The NCEAS dataset has been used to evaluate and compare various SDM methods based on presence-only data  (Elith et al., 2006; Phillips et al., 2009; Zbinden et al., 2024; Valavi et al., 2022) and it allows us to compare the performances of DeepMaxent with existing SDM methods (section below).
(Phillips et al., 2009) studied spatial sampling biases and found that the PO data in certain regions (AWT, CAN, and SWI) contain high levels of such biases. This is the data that we use to train all our models.
@@ 2.2.1 Dataset
Code Location Biological Group Species number Occurrences number
@@
AWT Australian wet tropics bird 20 3105 340
AWT Australian wet tropics plant 20 701 102
CAN Ontario, Canada bird 20 5063 14571
NSW New South Wales bates 10 187 570
NSW New South Wales birds 7 1781 1839
NSW New South Wales plants 29 680 5329
NSW New South Wales reptile 8 675 1008
NZ New Zealand plant 52 3088 19120
SA South America plant 30 2220 152
SWI Switzerland tree 30 35105 10013
@@ 2.2.2 Evaluation metrics
To directly compare our results to Phillips et al. (2009), Zbinden et al. (2024) and Valavi et al. (2022), we evaluated our method performances with the Area Under the ROC Curve (AUC) on the PA plots of the NCEAS dataset. The AUC is the empirical probability that a presence site has a higher model-predicted value than an absence site. In other words, it measures the model ability to distinguish between presence and absence classes based on its predicted scores. The NCEAS dataset being decomposed into regions and biological groups, for each region the average across biological groups of the AUCs and the general average of the latters across regions were calculated.
@@ 2.2.3 Implementation details
In this study, the first version of DeepMaxent method was implemented using Python and PyTorch framework and openly available at https://github.com/RYCKEWAERT/deepmaxent. The architecture was designed as a Multilayer Perceptron (MLP) residual neural networks (see Figure 1) with four fully connected layers interconnected by Rectified Linear Unit (ReLU) activation.
Cross-validation was performed to optimise hyperparameters, using spatial blocking based on geographic data (Valavi et al., 2019; Roberts et al., 2017). As suggested in Zbinden et al. (2024), the cross-validation is performed exclusively on PO data. Once cross-validation has been performed, the parameter values are set and are the same for all DeepMaxent models for all regions and target-Groups, so as to be comparable with other models. The final model is then calibrated with these hyperparameter values on the whole PO data and applied to the PA data.
@@ 2.2.4 Baseline losses
We implemented the Poisson regression loss (see Eq. 1) to test the effect of the density normalisation in DeepMaxent on the estimator quality. The Poisson loss is also equivalent to fitting the intensity of a gridded inhomogeneous Poisson point process (Renner and Warton (2013)) over our sites. Other commonly used loss functions in deep learning, and notably for SDMs, namely Cross-Entropy over species (CE, Deneu et al. (2021); Brun et al. (2024)) and Binary Cross-Entropy (BCE, Benkendorf and Hawkins (2020); Zbinden et al. (2024)) were implemented. All the baseline implementations were performed using the same architecture (2 hidden layers), optimiser (Adam), and hyperparameters (learning rate: $0.0002$, batch size: $250$) as the best performing DeepMaxent implementation, except for the weight decay $\tau$ which we kept to 00 for the baselines, as this regularisation did not improve any of them. Each loss function was evaluated both with and without TGB correction.
The Cross-Entropy over species loss (CE), $\mathcal{L}_{\mathrm{CE}}(\lambda,y)$ (equation 9), measures for each site the deviation between a predicted probability distribution across species and the associated empirical distribution based on the species observations in that site. In this case, the predicted probabilities are obtained by normalising the intensity values over the species $\lambda:=\{\lambda_{ij}\}_{i=1,j=1}^{K,N}$ (implemented with a softmax as for DeepMaxent):
$\mathcal{L}_{\mathrm{CE}}(\lambda,y)=-\frac{1}{K}\sum_{i=1}^{K}\sum_{j=1}^{N}\frac{y_{ij}}{\sum_{k=1}^{N}y_{ik}}\log\left(\frac{\lambda_{ij}}{\sum_{k=1}^{N}\lambda_{ik}}\right)$ (9)
The Binary Cross-Entropy loss, $\mathcal{L}_{\mathrm{BCE}}(\lambda,y_{\mathbb{b}})$, was implemented for the case where $y_{\mathbb{b}}\in\{0,1\}$ represents a binary variable, taking the value 1 for presence and 0 for absence. For binary classification, a sigmoid function is used to map the predicted logits $\lambda_{ij}$ to probabilities, making it a simpler case compared to multi-class classification (which uses the softmax function). The BCE loss is defined as:
$\mathcal{L}_{\mathrm{BCE}}(\lambda,y_{\mathbb{b}})=-\frac{1}{KN}\sum_{i=1}^{K}\sum_{j=1}^{N}\left(y_{\mathbb{b},ij}\log\sigma(\log(\lambda_{ij}))+(1-y_{\mathbb{b},ij})\log(1-\sigma(\log(\lambda_{ij})))\right),$ (10)
where $\sigma$ is the sigmoid function here applied to the linear predictor, or "logit", of each species $\log(\lambda_{ij})=\gamma_{j}^{T}g_{\theta}(x_{i})+b_{j}$.

### v4

@@ 2.1.3 Batched algorithm and partition function approximation in DeepMaxent
One of the challenges of adapting Maxent to a deep learning framework is the computation of the partition functions $\sum_{k}\lambda_{kj}$ for species $j$, corresponding to the denominator in Equation 2, which normalises the predicted intensity $\lambda_{j}(\mathbf{x})$ over the $K$ sites. This may become problematic when the number of observed sites increases, which often happens when considering a large geographic domain or a finer spatial resolution. In this case, the number of terms $K$ in the partition function becomes large, making it challenging to compute it exactly. To address this challenge efficiently, we leverage standard stochastic optimisation techniques widely used in deep learning (e.g., mini-batch Stochastic Gradient Descent (Bottou and Laboratories, 1991)), with the specificity that we approximate the normalisation within mini-batches. Specifically, we compute the loss function on a small random subset of sites $B\subset\{1,\dots,K\}$, called mini-batch, hence normalising the intensities within the mini-batch. The mini-batch-wise loss is then written as follows:
@@
$\mathcal{L}(\tilde{\boldsymbol{\Lambda}}_{i\in B},\tilde{\mathbf{Y}}_{i\in B})=-\frac{1}{|B|N}\sum_{i\in B}\sum_{j=1}^{N}\frac{y_{ij}}{\sum_{i\in B}y_{ij}+\epsilon}\log\left(\frac{\lambda_{j}(\mathbf{x}_{i})}{\sum_{i\in B}\lambda_{j}(\mathbf{x}_{i})}\right).$ (4)
@@
Note that the denominator $\sum_{i\in B}\lambda_{j}(\mathbf{x}_{i})$ is strictly positive under typical model assumptions where the Poisson intensity functions $\lambda_{j}(\mathbf{x}_{i})$ are positive. However, the denominator $\sum_{i\in B}y_{ij}$, representing the count sum for class $j$ in the mini-batch, may be zero if no samples of class $j$ are present in the mini-batch. In practice, this requires either ensuring that each mini-batch contains at least one example of every class to avoid division by zero, or the addition of a small, but positive, $\epsilon$. This mini-batch-wise loss makes the model optimisation computationally feasible by computing it one mini-batch at a time and avoiding computing the full partition functions for every iteration, and thus allows it to be scalable to large domains and trained efficiently with any optimiser based on random mini-batches, such as the commonly used Adam (Kingma, 2014). However, it must be noted that the batch-wise loss is not a simple approximation of the full loss as, for instance, the normalised occurrences tend to have a larger value for smaller mini-batches. Nevertheless, we provide the mathematical guarantee that, for any mini-batch size $n$ ($1<n<K$), minimizing the model loss on all mini-batches also minimizes the full loss (see Appendix A.2). This suggests that our final estimator should be close to the global minimizer of the full loss, even though there is not guarantee to obtain the latter due to the non-convexity induced by the non-linear feature extractor. Similarly to supervised contrastive methods (Khosla et al., 2020), the intensity predictions could become more specific, and more concentrated around the occurrences, as the mini-batch size increases. A small mini-batch size could, on the other hand, result in smoother species intensities. Therefore, we tested the impact of the mini-batch size during training on the final predictions of DeepMaxent. Not that, given the expression of $\lambda_{j}(\mathbf{x}_{i})$ in Equation 3, the normalised intensities across the mini-batch $B$ used in Equation 4 are given by $\exp(\gamma_{j}^{\top}g_{\theta}(\mathbf{x}_{i}))/\sum_{k\in B}\exp(\gamma_{j}^{\top}g_{\theta}(\mathbf{x}_{k}))$. This is a particular case of the normalisation function commonly referred to as softmax, applied to the logits $\gamma_{j}^{\top}g_{\theta}(\mathbf{x}_{k})$ over the mini-batch $B$.
@@ 2.1.4 Spatial sampling bias correction with Target-Group Background correction
When occurrence concentration is biased by spatial variations in sampling effort, a popular SDM correction approach is the Target-Group Background (TGB) method (Phillips et al., 2009), which was initially proposed to correct sampling bias in Maxent. The method basically approximates the spatial sampling effort through the distribution of occurrences of a Target-Group (TG) of species, providing background points to Maxent for each site where TG species were reported. In other words, TGB restricts the study domain to the sites with at least an evidence of sampling effort (one observation), which reduces the problem of false absences. The strategy is expected to work when TG species are reported jointly with the focal species (e.g. the TG is a biological group targeted by the same citizen science program).
In DeepMaxent, which models multiple species simultaneously, the TGB strategy emerges implicitly by considering all samples within a mini-batch in the computation of the partition function. This makes the implementation of TGB particularly efficient, as the calculation of the intensity prediction for each species is recycled to be used for the TGB-enhanced partition function.
@@ 2.1.5 L2-regularisation implementation in DeepMaxent
Maxent makes use of L1 penalisation, on the species weights $\gamma_{j}$ that model the relation between the features and the density prediction. The L1 term, known as LASSO penalty, encourages $\gamma_{j}$ to become sparse, thus selecting a subset of features. The L1 regularisation is important in Maxent due to a number of features that grows more than quadratically with the number of environmental variables. In DeepMaxent, the latent features are learnt to maximise prediction performances and can be kept to a fixed dimensionality, removing the need for feature selection (Goodfellow et al., 2016). For DeepMaxent, we employ L2 regularisation, i.e. a penalty on the Euclidean norm of the $\gamma_{j}$ and the rest of the model weights, which encourages small but non-zero weights, which tends to induce a smoothing of the estimated species intensities. The intercept term $\mathbf{b}$, on the other hand, is not penalized, since it controls only the baseline level of the intensity function per species and does not affect the effective capacity of the model. The total loss function thus becomes:
@@ 2.2 Evaluation of model performance
2.2.1 Datasets
For our experiments we used two openly available datasets: (i) one from the National Centre for Ecological Analysis and Synthesis (NCEAS) (Elith et al., 2020) and (ii) GeoPlant (Picek et al., 2025).
The NCEAS dataset
This dataset includes 52,605 PO records (for SDM training) and presence-absence (PA, for SDM evaluation) data from six global regions: Australian Wet Tropics (AWT), Canada (CAN), New South Wales (NSW), New Zealand (NZ), South America (SA) and Switzerland (SWI) (Elith et al., 2020). Each regions is associated with a specific set of species, and sometimes from several biological groups (see Table 1), with a total of 226 anonymous species. The dataset provides specific environmental variables with specific spatial resolution for each region, including climatic, soil or location variables (see more details in Elith et al. (2020)). The total area of all regions is of 13,607,500 km2 (Table 1), so that there is $0.004$ PO records per km2 on average when pooling all species and regions. However, the spatial concentration of records is extremely heterogeneous across regions and taxonomic groups, varying from $0.0002$ records/km2 in South America (SA) to $0.887$ records/km2 in Switzerland (SWI), as computed from Table 1.
Various SDM methods have been evaluated using this dataset (Elith et al., 2006; Phillips et al., 2009; Zbinden et al., 2024; Valavi et al., 2022), which allows for a direct comparison of DeepMaxent’s performance to many state-of-the-art SDM methods. Phillips et al. (2009) studied spatial sampling biases and found that the PO data in certain regions (AWT, CAN, and SWI) contained high levels of such biases, making this a good benchmark to assess model robustness.
@@ 2.2.1 Datasets
Code Location Biological Group Species number Occurrences number Area (‘000 km2)
@@
AWT Australian wet tropics bird 20 3105 340 24
AWT Australian wet tropics plant 20 701 102 24
CAN Ontario, Canada bird 20 5063 14571 979.3
NSW New South Wales bates 10 187 570 76.2
NSW New South Wales birds 7 1781 1839 76.2
NSW New South Wales plants 29 680 5329 76.2
NSW New South Wales reptile 8 675 1008 76.2
NZ New Zealand plant 52 3088 19120 265.4
SA South America plant 30 2220 152 12223.2
SWI Switzerland tree 30 35105 10013 39.6
The GeoPlant dataset
Described by Picek et al. (2025), we use this dataset to evaluate DeepMaxent under a different data regime, with 100 times more PO records, much more concentrated in space, and 40 times more species covered, than NCEAS. Indeed, GeoPlant contains 5,079,797 presence-only observations of 9,709 plant species (i.e. almost half of Europe’s flora) from 13 selected datasets of the Global Biodiversity Information Facility (GBIF, www.gbif.org), to be used for model training. These records cover 38 European countries spanning a total area of about 5,914,500 km2, with $0.859$ records per km2 on average. This spatial concentration in Geoplant is thus 200 times higher than the average of NCEAS, and comparable to the one of the SWI region, but for an area 150 times larger. Additionally, 88,987 presence-absence records from the European Vegetation Archive (EVA) were used for model evaluation.
Within GeoPlant, two types of input data were considered in this work for training SDMs. The first configuration (Bioclim-GeoPlant) used bioclimatic variables aggregated to a 10 km resolution, enabling direct comparison with MaxEnt. The second configuration (LandSat-GeoPlant) focused on applying deep learning-based approaches to multi-band time series of satellite data. Only methods that include a neural network feature extractor were considered here. We used Landsat-based covariates at 30 m resolution, derived from the seasonally aggregated and gap-filled GLAD analysis-ready dataset (Potapov et al., 2020), and accessed via the EcoDataCube platform (Witjes et al., 2022). These data covered six spectral bands: Red (R), Green (G), magenta (B), Near Infrared (NIR), Shortwave Infrared 1 (SWIR1), and Shortwave Infrared 2 (SWIR2).
@@ 2.2.2 Evaluation metrics
To directly compare our results to Phillips et al. (2009), Zbinden et al. (2024) and Valavi et al. (2022), we evaluated our method performances with the Area Under the ROC Curve (AUC) computed for each species across the PA plots, none of which were used in model training. The AUC is the empirical probability that a presence site has a higher model-predicted value than an absence site. In other words, it measures the model ability to distinguish between presence and absence classes based on its predicted scores. The NCEAS dataset being decomposed into regions and biological groups, we first averaged each species-wise AUC per biological group and then per region (across groups), and then took the average over regions as our general performance metric.
@@ 2.2.3 Implementation details
NCEAS Dataset
For the NCEAS dataset, the feature extractor $g_{\theta}$ was implemented as a multilayer perceptron (MLP) with rectifier linear unit (ReLU) non-linearities and skip connection between hidden layers (see Figure 5) in order to mitigate potential vanishing gradient issues when exploring deeper architectures (He et al., 2016). The use of an MLP as a feature extractor was justified both by the structure of the input data (vectors of environmental variables for each site) and as a standard reference in deep learning-based multi-species SDMs  (Zbinden et al., 2024; Kellenberger et al., 2024; Hu et al., 2025).
In this study, the parameters related to the neural network architecture and the optimiser, referred to as hyper-parameters, were selected through a cross-validation procedure. The search included the number of hidden layers, learning rate, mini-batch size and weight decay, and was performed using spatially blocked folds based on geographic data (Valavi et al., 2019; Roberts et al., 2017). As suggested in Zbinden et al. (2024), the cross-validation was performed using PO data (see details in appendix C). All models were trained for 100 epochs. Once cross-validation has been performed, the hyper-parameter values were chosen to be the same for all regions and biological-groups. These hyper-parameters were: 2 hidden layers, Adam as optimiser, a learning rate of $0.0002$, and a both the mini-batch size and hidden layer size of $250$. Each loss function was evaluated both with and without TGB correction. Among the hyper-parameters tested, these values consistently yielded the best results for all loss functions (see Annex C). Regarding weight decay $\tau$, we did not observe any performance improvement across the tested loss functions, except in the case of DeepMaxent, where it led to better results. The final model was then calibrated with these hyper-parameter values on the whole PO data and applied to the PA data. To account for variability arising from model initialization, each model was trained and evaluated across ten different random seeds, following the same procedure as used in Zbinden et al. (2024).
For the NCEAS dataset, although DeepMaxent implicitly implements Target Group Background (TGB) correction, we additionally added random background points in locations with no species observations to verify that adding such points would not contribute to improve model performance. Background points were sampled uniformly from the raster data, producing a dataset ten times larger than the presence-only (PO) occurrences. This procedure ensured a consistent representation of environmental conditions across the study area and maintained a percentage-based approach to guarantee comparability across datasets (e.g., SWI or CAN).
GeoPlant Datasets
For the Bioclim-GeoPlant dataset, we used the same MLP architecture as for NCEAS. This choice was again driven by the structure of the data, consisting of 19 bioclimatic variables. For the LandSat-GeoPlant dataset, each occurrence in our dataset was represented as a multidimensional data cube of 6 spectral bands × 4 seasons × 21 years, which serves as input to the model (see Figure 24). As a feature extractor we used the adapted ResNet-18 model proposed in Picek et al. (2025). The output features from the ResNet-18 were processed through two fully connected layers, resulting in the final predictions. The number of epochs was fixed at 20 throughout the training process and each hidden layer was composed of 250 neurons.
For both GeoPlant datasets, cross-validation was not performed due to computational constraints. Instead, a validation set was created by randomly splitting the presence-only (PO) data to select the best model. This approach ensured that model performance was assessed without excessive computational overhead. The hyper-parameters used for training were adapted from Picek et al. (2025), maintaining consistency with previously established configurations.
@@ 2.2.4 Baseline losses
We implemented the Poisson regression loss (see Equation 1), using the same neural network architecture as in DeepMaxent to model $\lambda_{j}$, to test the effect of the density normalisation in DeepMaxent on the estimator quality. Other commonly used loss functions in deep learning, and notably for SDMs, namely Cross-Entropy over species (CE, Deneu et al. (2021); Brun et al. (2024)) and Binary Cross-Entropy (BCE, Benkendorf and Hawkins (2020); Zbinden et al. (2024)) were implemented. The Cross-Entropy loss (CE) over species, $\mathcal{L}_{\mathrm{CE}}(\boldsymbol{\Lambda},\mathbf{Y})$ (Equation 6), measures for each site the deviation between a predicted probability distribution across species and the associated empirical distribution based on the species observations in that site. In this case, the predicted probabilities are obtained by normalising the intensity values over the species $\boldsymbol{\Lambda}:=\{\lambda_{ij}\}_{i=1,j=1}^{K,N}$, at each site separately, implemented using the softmax function:
$\mathcal{L}_{\mathrm{CE}}(\boldsymbol{\Lambda},\mathbf{Y})=-\frac{1}{K}\sum_{i=1}^{K}\sum_{j=1}^{N}\frac{y_{ij}}{\sum_{k=1}^{N}y_{ik}}\log\left(\frac{\lambda_{ij}}{\sum_{k=1}^{N}\lambda_{ik}}\right)$ (6)
The Binary Cross-Entropy (BCE) loss, $\mathcal{L}_{\mathrm{BCE}}(\boldsymbol{\Lambda},\mathbf{Y}_{\mathbb{b}})$, was implemented for the case where $\mathbf{Y}_{\mathbb{b}}\in\{0,1\}$ is treated binary variable of $\mathbf{Y}$, taking the value 1 if the species was observed at least once in the pixel, and 0 otherwise. Unlike the CE case, where the probabilities across species are required to sum to 1 in each site, the predicted probability of a species learned using BCE does not directly restrict the probabilities of other species. In this setting, the softmax function reduces to a sigmoid function. The BCE loss is defined as:
$\mathcal{L}_{\mathrm{BCE}}(\boldsymbol{\Lambda},\mathbf{Y}_{\mathbb{b}})=-\frac{1}{KN}\sum_{i=1}^{K}\sum_{j=1}^{N}\left(y_{\mathbb{b},ij}\log\sigma(\log(\lambda_{ij}))+(1-y_{\mathbb{b},ij})\log(1-\sigma(\log(\lambda_{ij})))\right),$ (7)
where $\sigma$ is the logistic sigmoid function here applied to the linear predictor, or "logit", of each species $\log(\lambda_{ij})=\gamma_{j}^{\top}g_{\theta}(\mathbf{x}_{i})+b_{j}$.